# Dual-scale Enhanced and Cross-generative Consistency Learning for Semi-supervised Polyp Segmentation

### v1

@@ III-D Dual-scale Complementary Fusion Module
[…]
@@
Subsequently, the fused features $h_{i}^{dcf}$ will be passed into the fused decoder. Specifically, $h_{5}^{dcf}$ is first passed through a transpose convolutional layer to have the same resolution with $h_{4}^{dcf}$. Then, the two features are concatenated and fed into two convolutional layers. Repeating the above process, we will obtain the final segmentation maps $S_{l}^{f}$ for labeled data. Therefore, the supervised loss in Eq. (4) can be reformulated as follows:
$\begin{split}\mathcal{L}_{S}=&\mathcal{L}_{sup}(S_{l},Y_{l})+\\ &\mathcal{L}_{sup}(S_{l}^{d},Down(Y_{l}))+\mathcal{L}_{sup}(S_{l}^{f},Y_{l}).\end{split}$ (6)
[…]
Figure 4: Illustration of (a) the architecture of the fused decoder (i.e., ${D}_{s}^{f}$ and ${D}_{t}^{f}$ in Fig. 1) and (b) the proposed dual-scale complementary fusion module.
@@ III-E Scale-aware Perturbation and Cross-generative Consistency
Following previous works [35, 33, 27, 42, 43], to improve the robustness of the MT model, $X_{u}$ and its perturbed version $X_{u}^{p}$ are passed through both student and teacher models, respectively, to encourage the outputs to be closed. In this study, we design a scale-aware perturbation consistency. Specifically, we impose the perturbation consistency on different scale images, i.e., $S_{u}^{p}=D_{t1}(E_{t}(\mathcal{P}(X_{u})))$ and $S_{u}^{dp}=D_{t2}(E_{t}(\mathcal{P}(Down(X_{u}))))$, where a random gamma operation is adopted as the perturbation operation (i.e., $\mathcal{P}$). Meanwhile, we also obtain ${S}_{u}^{fp}$ that its constraint is implemented by the MT framework, due to the particularity of ${S}_{u}^{fp}$. Thus, the scale-aware perturbation consistency scheme can be optimized by minimizing $\mathcal{L}_{SPC}$, which is defined by
$\mathcal{L}_{SPC}=\mathcal{L}_{mse}(S_{u},S_{u}^{p})+\mathcal{L}_{mse}(S_{u}^{d},S_{u}^{dp})+\mathcal{L}_{mse}(S_{u}^{f},S_{u}^{fp}).$ (7)
@@
Further, to utilize the rich information contained in different feature maps, we propose a perturbation-based cross-generative consistency constraint to enhance the feature mining ability of our segmentation network. Therefore, we design two generative networks $G_{1}$ and $G_{2}$ (The detailed structure is shown in Fig. 5) for image reconstruction. The segmentation maps $S_{u}^{f}$ and $S_{u}^{fp}$ obtained from the aggregation of the two scale information are fed into $G_{2}$ and $G_{1}$, respectively, and then we can obtain two reconstructed images $\hat{X}_{u}=G_{1}(S_{u}^{fp})$ and $\hat{X}_{u}^{p}=G_{2}(S_{u}^{f})$. Finally, we form a cross-generative consistency loss ($\mathcal{L}_{CC}$), which is given by
$\mathcal{L}_{CC}=\mathcal{L}_{mse}(\hat{X}_{u},X_{u})+\mathcal{L}_{mse}(\hat{X}_{u}^{p},X_{u}^{p}).$ (8)
It is worth noting that $\mathcal{L}_{CC}$ can enforce $\hat{X}_{u}$ and $X_{u}$ to be consistent and $\hat{X}_{u}^{p}$ and $X_{u}^{p}$ to be closed. As a result, such cross-generation can further improve the robustness of our MT-based model to learn powerful features and fully harness the knowledge from unlabeled data to boost the proposed segmentation model.
@@
CVC-ClinicDB Baseline 0.767 0.689 0.756 0.847 0.024 0.812 0.751 0.806 0.873 0.021
MT [15] 0.785 0.711 0.763 0.855 0.025 0.839 0.781 0.829 0.892 0.020
DAN [34] 0.789 0.718 0.779 0.862 0.024 0.838 0.781 0.835 0.890 0.018
UA-MT [10] 0.798 0.729 0.782 0.869 0.022 0.831 0.762 0.817 0.879 0.026
URPC [16] 0.794 0.723 0.783 0.864 0.022 0.829 0.767 0.825 0.880 0.022
Duo-SegNet [32] 0.765 0.685 0.745 0.840 0.037 0.755 0.678 0.739 0.832 0.038
@@
BCP [44] 0.835 0.769 0.828 0.887 0.018 0.846 0.782 0.847 0.894 0.020
DEC-Seg (ours) 0.830 0.772 0.836 0.888 0.014 0.852 0.798 0.853 0.897 0.019
Kvasir Baseline 0.786 0.701 0.756 0.831 0.059 0.843 0.782 0.829 0.877 0.037
MT [15] 0.833 0.755 0.818 0.858 0.051 0.868 0.799 0.853 0.881 0.037
DAN [34] 0.836 0.758 0.820 0.858 0.049 0.867 0.803 0.851 0.886 0.035
UA-MT [10] 0.824 0.753 0.808 0.859 0.051 0.858 0.787 0.838 0.881 0.039
URPC [16] 0.821 0.740 0.803 0.853 0.050 0.854 0.789 0.843 0.879 0.037
Duo-SegNet [32] 0.795 0.711 0.785 0.838 0.061 0.851 0.772 0.843 0.878 0.043
@@
BCP [44] 0.854 0.789 0.850 0.876 0.042 0.886 0.820 0.876 0.898 0.029
DEC-Seg (ours) 0.862 0.794 0.853 0.880 0.040 0.882 0.822 0.871 0.899 0.033
III-F Loss Function and Optimization
Inspired by [45], the supervised loss for labeled data is formulated by $\mathcal{L}_{sup}=\mathcal{L}_{IoU}^{w}+\mathcal{L}_{BCE}^{w}$, where $\mathcal{L}_{IoU}^{w}$ and $\mathcal{L}_{BCE}^{w}$ denote the weighted IoU loss and binary cross-entropy (BCE) loss, respectively. Finally, the total loss can be given as follows:
$\mathcal{L}_{total}=\mathcal{L}_{S}+\mathcal{L}_{SPC}+\alpha\mathcal{L}_{SC}+\beta\mathcal{L}_{CC},$ (9)
where the hyper-parameter $\alpha$ is set to $1.5$ and $\beta$ is set to $3$. In addition, the parameters $\theta^{{}^{\prime}}$ of the teacher network can be updated based on the parameters $\theta$ of the student network using an exponential moving average (EMA) strategy. Thus, at the $t$-th training step, the parameters $\theta^{{}^{\prime}}$ can be updated by $\theta^{{}^{\prime}}_{t}=\gamma\theta^{{}^{\prime}}_{t-1}+(1-\gamma)\theta_{t}$, where $\gamma$ is used to control the EMA decay, and it is set to $0.99$.
IV EXPERIMENTS AND RESULTS
IV-A Experimental Setup
IV-A1 Datasets and Training Settings
We carry out the comparison experiments on five public datasets, i.e., ETIS [46], CVC-ClinicDB [47], EndoScene-CVC300 [48], CVC-ColonDB [49], and Kvasir [50]. Following the setting in [2], $1450$ images are randomly selected from the two datasets (Kvasir and CVC-ClinicDB) for the training set, and the remaining images from the two datasets and the other three datasets (EndoScene-CVC300, ETIS, and CVC-ColonDB) to form the testing set. Besides, $10\%$ ($145$ images) or $30\%$ ($435$ images) from the training set are used as the labeled data, while the remaining images are adopted as the unlabeled data.
IV-A2 Implementation Details and Evaluation Metrics
The proposed segmentation framework is implemented in PyTorch and trained on one NVIDIA GeForce RTX3090 GPU with an Adam optimizer. All inputs are uniformly rescaled to $352\times 352$. Our model converges over $10000$ iterations with a batch size of $8$ including 4 labeled samples and 4 unlabeled samples, and the learning rate is set to $1e$-$4$.
To evaluate the effectiveness, we employ five commonly adopted metrics [2, 3], namely mean Dice (mDice), mean IoU (mIoU), $F_{\beta}^{w}$, $S_{\alpha}$, and mean absolute error (MAE).
TABLE II: Quantitative results on three unseen datasets (EndoScene-CVC300, CVC-ColonDB, and ETIS) with $10\%$ or $30\%$ labeled data.
@@ IV-A2 Implementation Details and Evaluation Metrics
EndoScene-CVC300 Baseline 0.800 0.708 0.784 0.861 0.013 0.827 0.760 0.812 0.895 0.008
MT [15] 0.858 0.779 0.837 0.910 0.010 0.873 0.796 0.845 0.916 0.009
DAN [34] 0.820 0.733 0.797 0.885 0.014 0.874 0.800 0.847 0.917 0.008
UA-MT [10] 0.859 0.780 0.829 0.910 0.010 0.837 0.754 0.795 0.895 0.012
URPC [16] 0.841 0.753 0.812 0.896 0.012 0.872 0.800 0.859 0.920 0.007
Duo-SegNet [32] 0.841 0.761 0.811 0.894 0.011 0.843 0.758 0.803 0.902 0.012
CLCC [17] 0.748 0.674 0.724 0.845 0.015 0.826 0.755 0.811 0.891 0.008
SLC-Net [18] 0.856 0.777 0.827 0.907 0.012 0.869 0.796 0.849 0.915 0.010
MC-Net+ [37] 0.783 0.711 0.755 0.872 0.014 0.845 0.783 0.844 0.891 0.020
CDMA [38] 0.791 0.685 0.765 0.846 0.013 0.803 0.728 0.802 0.863 0.024
BCP [44] 0.856 0.770 0.835 0.901 0.010 0.846 0.782 0.847 0.894 0.020
DEC-Seg (ours) 0.860 0.781 0.831 0.914 0.011 0.895 0.824 0.875 0.927 0.006
CVC-ColonDB Baseline 0.531 0.453 0.526 0.712 0.052 0.551 0.489 0.543 0.732 0.049
MT [15] 0.610 0.531 0.597 0.759 0.044 0.668 0.587 0.652 0.788 0.041
DAN [34] 0.599 0.515 0.593 0.748 0.049 0.629 0.551 0.613 0.766 0.043
UA-MT [10] 0.583 0.507 0.576 0.745 0.050 0.666 0.587 0.641 0.785 0.048
URPC [16] 0.587 0.507 0.574 0.745 0.048 0.621 0.551 0.612 0.768 0.045
Duo-SegNet [32] 0.644 0.550 0.613 0.763 0.058 0.658 0.569 0.627 0.773 0.057
@@
BCP [44] 0.642 0.563 0.635 0.773 0.046 0.656 0.580 0.651 0.778 0.046
DEC-Seg (ours) 0.650 0.575 0.639 0.782 0.042 0.691 0.616 0.678 0.798 0.042
ETIS Baseline 0.458 0.392 0.433 0.693 0.024 0.486 0.432 0.469 0.715 0.021
MT [15] 0.468 0.395 0.434 0.694 0.036 0.595 0.519 0.552 0.761 0.035
DAN [34] 0.473 0.395 0.440 0.700 0.031 0.528 0.452 0.485 0.729 0.045
UA-MT [10] 0.469 0.399 0.444 0.695 0.027 0.603 0.515 0.547 0.761 0.032
URPC [16] 0.459 0.391 0.436 0.691 0.028 0.601 0.523 0.572 0.765 0.023
Duo-SegNet [32] 0.450 0.382 0.407 0.669 0.061 0.578 0.500 0.531 0.750 0.049
@@
BCP [44] 0.571 0.502 0.550 0.756 0.023 0.557 0.497 0.541 0.754 0.020
DEC-Seg (ours) 0.615 0.540 0.582 0.775 0.019 0.683 0.609 0.658 0.814 0.014
IV-B Comparison with State-of-the-arts

### v2

@@ III-D Dual-scale Complementary Fusion Module
The two independent decoders, as designed in Section III-C, are specialized for processing information at two distinct scales. Throughout network training, these decoders acquire unique features based on the input image scale, with each emphasizing different aspects. Despite this distinction, the features from both scales pertain to the same unlabeled image, indicating the opportunity to combine and extract features from these two scale features further, potentially resulting in an enhanced segmentation map. Consequently, to fully leverage information from different scales, in comparison to the scale-specific decoder, we also introduce the scale-fused decoder (Df in Fig. 1) to produce more reliable predictions as the final segmentation maps.
As shown in Fig. 1, the features from $D_{1}$ and $D_{2}$ are fused and then fed into the scale-fused decoder. The feature flow of the scale-fused decoder is shown in Fig. 4(a), and the structure of the scale-fused decoder is basically the same as that of the scale-specific decoder. To achieve the fuse operation, we present a dual-scale complementary fusion (DCF) module to fuse the features from the two scale-specific decoders. This module replaces the features transmitted from the encoder through the skip operation. The proposed DCF module aims to fuse the multi-scale features and enable the features from the original scale and downsampled scale to complement each other (as shown in Fig. 4(b)). Taking the original scale feature as an example, we first conduct a $3\times 3$ convolution on $Up(h_{i}^{2})$ and Sigmoid activation function to obtain a scale-aware weight $W_{2}=\sigma(\mathcal{B}_{conv3\times 3}(Up(h_{i}^{2})))$. Then, we multiply $W_{2}$ by $\mathcal{B}_{conv3\times 3}(Up(h_{i}^{2}))$, and the resultant feature is the complementary information (i.e., $h_{com}^{2}$) which is encouraged to learn the information required by the original scale. Finally, we add the complementary feature obtained at the downsampled scale to the original scale and then smooth the feature by a $3\times 3$ convolution, thus we can obtain $h_{rec}^{1}=\mathcal{B}_{conv3\times 3}(\mathcal{B}_{conv3\times 3}(h_{i}^{1})+\mathcal{B}_{conv3\times 3}(Up(h_{i}^{2}))*W_{2})$. Similarly, we can obtain the complementary-enhanced downsampled feature $h_{rec}^{2}$. Through a cross-scale complementary-enhanced process, we learn more rich multi-scale feature representations. Further, two features $h_{rec}^{1}$ and $h_{rec}^{2}$ are concatenated and then proceeded by a $1\times 1$ convolution to two channels and a Softmax function to obtain two weight maps $\alpha_{1}$ and $\alpha_{2}$, where $\alpha_{1}+\alpha_{2}=1$. Finally, we obtain the multi-scale fused feature by
@@
Subsequently, the fused features $h_{i}^{dcf}$ will concatenate with the features from the previous layer and continue to propagate forward. Specifically, $h_{5}^{dcf}$ is first passed through two convolutional layers and a transpose convolutional layer to have the same resolution with $h_{4}^{dcf}$. Then, the two features are concatenated and fed into the next two convolutional layers. Repeating the above process, we will obtain the final segmentation maps $S^{f}$. It is important to highlight that the scale-fused decoder plays a vital role in fully integrating multi-scale information, consequently enhancing the segmentation performance. Consequently, during the inference stage, the segmentation predictions obtained from the scale-fused decoder are taken as the final segmentation results.
@@ III-E Scale-aware Perturbation and Cross-generative Consistency
To constrain the predictions of unlabeled data on the scale-fused decoder, we follow previous works [25, 40, 41], and introduce scale-aware perturbation consistency to encourage the outputs between the original version and perturbed version to be closed. Specifically, we impose the perturbation consistency on different scale images, i.e., $S_{p}^{1}=D_{1}(E(\mathcal{P}(X_{u})))$, $S_{p}^{2}=D_{2}(E(\mathcal{P}(X_{u}^{2})))$, and $S_{p}^{f}$, where a random color jitter operation is adopted as the perturbation operation (i.e., $\mathcal{P}$). Thus, $\mathcal{L}_{SC}^{p}(S^{1}_{p},S^{2}_{p})$ is obtained and the scale-aware perturbation consistency scheme can be optimized by minimizing $\mathcal{L}_{SPC}$, which is defined by
$\mathcal{L}_{SPC}=\mathcal{L}_{mse}(S_{u}^{1},S_{p}^{1})+\mathcal{L}_{mse}(S_{u}^{2},S_{p}^{2})+\mathcal{L}_{mse}(S_{u}^{f},S_{p}^{f}).$ (5)
@@
Further, to utilize the rich information contained in different feature maps, we propose a perturbation-based cross-generative consistency constraint to enhance the feature mining ability of our segmentation network with information reconstruction and to correct the cognitive bias on a single input version with cross-generation. Therefore, we design two generative networks $G_{1}$ and $G_{2}$ (The detailed structure is shown in Fig. 5) for image reconstruction. For the generative network to accept richer information, produce better-reconstructed maps, and deliver more valuable feedback to the segmentation network, the logit predictions $z_{u}^{f}$ and $z_{p}^{f}$ obtained from the scale-fused decoder are fed into $G_{2}$ and $G_{1}$, respectively, and then we can obtain two reconstructed images $\hat{X}_{p}=G_{1}(z_{p}^{f})$ and $\hat{X}_{u}=G_{2}(z_{u}^{f})$. Finally, we form a cross-generative consistency loss ($\mathcal{L}_{CC}$), which is given by
$\mathcal{L}_{CC}=\mathcal{L}_{mse}(\hat{X}_{p},X_{u})+\mathcal{L}_{mse}(\hat{X}_{u},X_{p}).$ (6)
It is worth noting that $\mathcal{L}_{CC}$ can enforce $\hat{X}_{p}$ and $X_{u}$ to be consistent and $\hat{X}_{u}$ and $X_{p}$ to be closed. Using the input image as a learning target provides unlabeled data with an authentic constraint containing structural and semantic information. This empowers the segmentation network to prioritize the structural and semantic details that are pertinent to category information during the generation of segmentation maps. The cross-generation process compels the network to investigate inconsistent predictions caused by color variations, thereby rectifying the network’s cognitive biases toward a single input and ensuring the accuracy of predictions for both the original input and the perturbed input from different perspectives. Consequently, such cross-generation can further enhance the robustness of our model, enabling it to learn powerful features and effectively leverage the knowledge from unlabeled data to enhance the proposed segmentation model.
III-F Overall Loss Function
In this study, the supervised loss for labeled data is the sum of the loss on the prediction maps of the designed three decoders and its corresponding ground truth, and it is formulated by
$\begin{split}\mathcal{L}_{S}=\mathcal{L}_{sup}(S_{l}^{1},Y_{l})+\mathcal{L}_{sup}(S_{l}^{2},Down(Y_{l}))+\mathcal{L}_{sup}(S_{l}^{f},Y_{l}),\end{split}$ (7)
where $\mathcal{L}_{sup}=(\mathcal{L}_{CE}+\mathcal{L}_{Dice})/2$, $\mathcal{L}_{CE}$ and $\mathcal{L}_{Dice}$ denote the cross-entropy (CE) loss and Dice loss, respectively. Finally, the total loss can be given as follows:
$\mathcal{L}_{total}=\mathcal{L}_{S}+\mathcal{L}_{SPC}+\mathcal{L}_{SC}+\mathcal{L}_{CC},$ (8)
where $\mathcal{L}_{SC}=\mathcal{L}_{SC}^{l}+\mathcal{L}_{SC}^{u}+\mathcal{L}_{SC}^{p}$ is the sum of scale-enhanced consistency loss in labeled images, unlabeled images, and perturbed images.
IV Experiments and Results
IV-A Datasets
Colonoscopy Datasets: We carry out the comparison experiments on four public datasets, i.e., CVC-ColonDB [42], Kvasir [43], CVC-ClinicDB [44], and ETIS [45]. Following the setting in [46], $1,450$ images are randomly selected from the two datasets (i.e., CVC-ClinicDB and Kvasir) for the training set, and the remaining images from the two datasets and the other two datasets (i.e., ETIS and CVC-ColonDB) to form the testing set. Besides, $10\%$ ($145$ images) or $30\%$ ($435$ images) from the training set are used as the labeled data, while the remaining images are adopted as the unlabeled data.
Brain Tumor Dataset: This dataset contains brain MR images together with manual FLAIR abnormality segmentation masks and it comes from 110 patients included in The Cancer Genome Atlas (TCGA) lower-grade glioma [47, 48]. In this study, we have $1,090$ images for the training set and 283 ones for the testing set. In addition, $10\%$ and $30\%$ images from the training set are adopted as the labeled data, respectively, while the remaining training images are taken as unlabeled data.
Skin Lesion Dataset: The dataset comes from ISIC-2018 challenge [49] and is used for skin lesion segmentation. It includes $2,594$ dermoscopic images for training and $1,000$ images for testing. In the same way, $10\%$ or $30\%$ images from the training set are used as the labeled data, while the remaining images are formed as the unlabeled set.
@@ IV-A Datasets
CVC-ClinicDB MT [10] 0.747 0.662 0.741 0.830 0.028 0.772 0.689 0.763 0.852 0.028
DAN [32] 0.755 0.674 0.733 0.842 0.027 0.775 0.696 0.772 0.848 0.027
UA-MT [11] 0.749 0.676 0.742 0.839 0.027 0.790 0.715 0.787 0.853 0.025
URPC [14] 0.769 0.696 0.760 0.854 0.021 0.777 0.709 0.785 0.853 0.026
@@
SCP-Net [50] 0.776 0.703 0.758 0.854 0.024 0.839 0.786 0.837 0.888 0.020
MCF [51] 0.779 0.718 0.767 0.860 0.021 0.823 0.766 0.819 0.882 0.019
DEC-Seg (ours) 0.836 0.774 0.834 0.886 0.019 0.859 0.804 0.860 0.900 0.015
Kvasir MT [10] 0.814 0.722 0.797 0.841 0.050 0.814 0.732 0.790 0.852 0.052
DAN [32] 0.808 0.723 0.786 0.839 0.059 0.841 0.760 0.828 0.866 0.043
UA-MT [11] 0.799 0.713 0.782 0.834 0.058 0.845 0.771 0.840 0.873 0.037
URPC [14] 0.811 0.728 0.796 0.842 0.057 0.849 0.778 0.846 0.874 0.040
@@
SCP-Net [50] 0.810 0.723 0.790 0.840 0.056 0.862 0.789 0.847 0.878 0.037
MCF [51] 0.822 0.751 0.816 0.852 0.051 0.856 0.794 0.847 0.882 0.039
DEC-Seg (ours) 0.859 0.787 0.853 0.877 0.044 0.893 0.830 0.886 0.902 0.032
IV-B Implementation Details and Evaluation Metrics
Implementation Details: The proposed segmentation framework is implemented in PyTorch and trained on one NVIDIA GeForce RTX3090 GPU. The input images from the polyp segmentation dataset and skin lesion segmentation dataset are uniformly rescaled to $352\times 352$, and the inputs of the brain MRI segmentation dataset are uniformly rescaled to $256\times 256$. Our model converges over $10,000$ iterations with a batch size of $6$ including 3 labeled samples and 3 unlabeled samples, and uses an SGD optimizer with an initial learning rate set to $1e$-$2$ and a poly learning rate strategy to update the learning rate.
Evaluation Metrics: To evaluate the effectiveness, we employ five commonly adopted metrics [46, 19], namely mean Dice (mDice), mean IoU (mIoU), $F_{\beta}^{w}$, $S_{\alpha}$, and mean absolute error (MAE).
TABLE II: Quantitative results on two unseen datasets (CVC-ColonDB and ETIS) with $10\%$ or $30\%$ labeled data.
@@ IV-B Implementation Details and Evaluation Metrics
CVC-ColonDB MT [10] 0.589 0.497 0.581 0.743 0.048 0.600 0.510 0.597 0.749 0.045
DAN [32] 0.620 0.517 0.604 0.756 0.046 0.633 0.546 0.630 0.767 0.045
UA-MT [11] 0.553 0.471 0.551 0.727 0.051 0.648 0.563 0.635 0.771 0.046
URPC [14] 0.556 0.480 0.549 0.732 0.047 0.598 0.519 0.601 0.754 0.045
@@
SCP-Net [50] 0.577 0.495 0.559 0.739 0.049 0.694 0.615 0.683 0.802 0.039
MCF [51] 0.566 0.498 0.559 0.742 0.045 0.621 0.550 0.614 0.768 0.042
DEC-Seg (ours) 0.648 0.565 0.630 0.771 0.046 0.721 0.640 0.709 0.814 0.035
ETIS MT [10] 0.356 0.288 0.337 0.636 0.035 0.495 0.420 0.462 0.706 0.039
DAN [32] 0.437 0.358 0.397 0.666 0.046 0.485 0.418 0.472 0.708 0.027
UA-MT [11] 0.459 0.393 0.441 0.693 0.029 0.541 0.473 0.523 0.738 0.024
URPC [14] 0.420 0.356 0.406 0.677 0.028 0.559 0.490 0.549 0.749 0.021
@@
SCP-Net [50] 0.394 0.323 0.355 0.656 0.045 0.571 0.500 0.543 0.747 0.023
MCF [51] 0.425 0.367 0.407 0.679 0.025 0.487 0.425 0.469 0.707 0.022
DEC-Seg (ours) 0.592 0.511 0.558 0.758 0.025 0.634 0.564 0.608 0.789 0.019
TABLE III: Quantitative results on Brain MRI dataset with $10\%$ or $30\%$ labeled data.
Methods 10% labeled 30% labeled
MT [10] 0.491 0.404 0.512 0.705 0.015 0.674 0.561 0.704 0.792 0.012
DAN [32] 0.534 0.429 0.564 0.719 0.015 0.703 0.582 0.737 0.803 0.011
UA-MT [11] 0.549 0.435 0.584 0.721 0.015 0.703 0.586 0.736 0.807 0.011
URPC [14] 0.547 0.448 0.571 0.729 0.014 0.653 0.548 0.677 0.785 0.012
CLCC [15] 0.595 0.505 0.615 0.760 0.012 0.728 0.615 0.754 0.824 0.010
SLC-Net [16] 0.483 0.400 0.506 0.704 0.015 0.702 0.591 0.714 0.807 0.010
MC-Net+ [36] 0.580 0.481 0.595 0.749 0.015 0.746 0.640 0.766 0.834 0.009
CDMA [37] 0.495 0.408 0.508 0.705 0.015 0.627 0.526 0.647 0.775 0.011
SCP-Net [50] 0.559 0.467 0.571 0.741 0.014 0.643 0.544 0.660 0.783 0.012
MCF [51] 0.589 0.491 0.623 0.750 0.013 0.704 0.612 0.722 0.819 0.010
DEC-Seg (ours) 0.626 0.533 0.639 0.777 0.012 0.767 0.671 0.783 0.850 0.008
TABLE IV: Quantitative results on ISIC-2018 dataset with $10\%$ or $30\%$ labeled data.
Methods 10% labeled 30% labeled
MT [10] 0.822 0.738 0.809 0.811 0.104 0.844 0.760 0.822 0.825 0.087
DAN [32] 0.831 0.744 0.817 0.815 0.097 0.842 0.757 0.825 0.824 0.091
UA-MT [11] 0.839 0.752 0.818 0.820 0.094 0.845 0.762 0.829 0.829 0.086
URPC [14] 0.847 0.763 0.829 0.831 0.089 0.853 0.772 0.835 0.880 0.082
CLCC [15] 0.842 0.756 0.815 0.820 0.088 0.845 0.762 0.816 0.821 0.092
SLC-Net [16] 0.843 0.754 0.818 0.822 0.091 0.846 0.752 0.811 0.821 0.085
MC-Net+ [36] 0.848 0.767 0.829 0.832 0.087 0.862 0.782 0.841 0.842 0.076
CDMA [37] 0.858 0.777 0.842 0.839 0.083 0.868 0.788 0.848 0.846 0.073
SCP-Net [50] 0.852 0.770 0.835 0.835 0.085 0.853 0.769 0.831 0.834 0.083
MCF [51] 0.856 0.773 0.837 0.836 0.083 0.861 0.778 0.839 0.840 0.077
DEC-Seg (ours) 0.867 0.796 0.858 0.851 0.078 0.875 0.801 0.862 0.856 0.069
IV-C Comparison with State-of-the-arts